Lesson: 2nd Law of Thermodynamics and Entropy




1. If you know that the change in entropy of a system where heat was added is 40 J/K, and that the temperature of the system is 200 K, what is the amount of heat added to the system?




2. If I add 30J of heat to a system so that the final temperature of the system is 26.85 °C, what is the change in entropy of the system?




3. Suppose 0.2 kg of ice at 0°C is in 0.5 kg water at 30°C. Determine the change in entropy of the ice as it melts at 0°C.

1
Expert's answer
2022-01-13T09:15:05-0500

"1.\space Q=\Delta S·T=8000\text{ J}.\\\space\\
2. \space \Delta S=\dfrac{Q}{T}=0.1\text{ J/K}\\\space\\
3. \space \Delta S=\dfrac{[4200·0.5(30-0)/334000]·334000}{273}=231\text{ J/K}."